OpenStudy (anonymous):

Linda sold 39 tickets for her choirs concert for a total of $101.30. If an adult ticket sold for $2.50 and a children's ticket sold for $2.70, how many of each kind of ticket did Linda sell? Please help!

OpenStudy (anonymous):

x = adults's ticket and y = children's ticket \[2.5x+2.7y = 39\] \[x+y = 101.30\] use matrix, or substitution, or elimination \[2.5x + 2.7(-x+101.3) = 39\]\[2.5x-2.7x+273.51 = 39\]\[-.2x=-234.51\]\[x \approx 1173\] \[1173+y = 101.3\]\[y = -1071.7\] ??? this doesn't sound right??

OpenStudy (anonymous):

ill see